# Sign problem in tensor network contraction

### v1

@@ III Monte Carlo based contraction
Fig. LABEL:fig:free-energy shows the average $\Delta f$ as a function of $\lambda$ for different bond dimensions $D$, both for Haar-random orthogonal and unitary tensors. We can identify three different regimes: First, for $\lambda\leq\eta_{0}/D$ (with $\eta_{0}\approx 1$), $\Delta f$ is constant, with a value $\Delta f=\log D+\mathrm{const.}$, as can be seen from the data collapse for all $D$ except $D=2$ in Fig. LABEL:fig:free-energyb (with the constant depending on the choice of the random ensemble). Thus, $e^{\Delta f}\sim D^{N}$, which is consistent with the expected worst-case $D^{N}$ scaling (corresponding to exact contraction). However, once outside this regime, the $\Delta f(\lambda)$ for different $D$ collapse more or less perfectly, with only minor deviations for $D=2$. $\Delta f(\lambda)$ then exhibits two different regimes: For $\eta_{0}/D\leq\lambda\lesssim 1$, we see a behavior $\Delta f\sim-\log\lambda$ (as expected from the collapse in Fig. LABEL:fig:free-energyb and the transition at constant $\lambda D$). For $\lambda\gg 1$, the scaling behavior depends on the chosen distribution—for Haar-orthogonal tensors, $\Delta f$ decays super-exponentially, while for Haar-unitary tensors, $\Delta f\sim\lambda^{2}/4$ (Fig. LABEL:fig:free-energyd,e). The latter behavior can be understood from the fact that for Haar-orthogonal tensors, only a Gaussian tail contributes to negative amplitudes, while for Haar-unitary tensors which are shifted by $\lambda\gg 1$ into the positive, the phase of the tensor still fluctuates by $\sim 1/\sqrt{\lambda}$ around zero. For comparison, we have also considered tensor networks with real or complex Gaussian entries (shown in Fig. LABEL:fig:free-energy with semi-transparent markers), which show the same behavior to very high accuracy; one difference, however, is the behavior for large $\lambda$ in the real case, as for Haar-orthogonal tensors, $\Delta f\equiv 0$ as soon as $\lambda\geq D^{2}$, and thus, the decay for large $\lambda$ must be more rapid (as can be seen in Fig. LABEL:fig:free-energyd for the $D=2$ data).
@@
From Fig. LABEL:fig:free-energy, we conclude that on the one hand, for very small $\lambda$, the severity of the sign problem depends on the bond dimension of the tensor network, and increases with larger $D$. For $1/D\lesssim\lambda\lesssim 1$, this transitions into a regime where there is still a sign problem, which however is independent of the bond dimension. Only for $\lambda\gtrsim 1$, the sign problem starts to disappear—in particular, in the orthogonal case, the superexponential decay of $\Delta f(\lambda)$ implies the absence of a sign problem as long as $N$ does not grow superexponentially with $\lambda$. The latter transition is indeed what one would have naively expected, given that around $\lambda\gtrsim 1$, the entries of the tensors start to be predominantly positive, with the probability of negative entries rapidly decreasing as $\lambda$ is increased.
@@ IV Entanglement based contraction
Figure 4: Contraction of a 2D tensor network. One proceeds by blocking columns (blue box in panel (a)). The resulting block $\psi$ is then approximated by a 1D tensor network, i.e., an MPS, with bond dimension $\chi$ (panel (b)). In order to keep the truncation error in all iterations small, it is necessary that the correlations between the upper and lower part of $\psi$ are bounded, i.e., the singular value decomposition in the $A$ vs. $B$ partition has a small tail (shown in panel (c) across the central cut, where we expect most correlations.) This weight can be assessed by whether the entropy across the cut scales with the width $W$. By choosing $H/2$ sufficiently larger than $W$, we ensure that solely $W$ bounds the possible correlations. Concretely, we choose $H/2=2W$; we have found that further increasing $H$ does not lead to an increase in correlations.
@@ IV.1 Numerical results
Leaving the surprising transition point aside, the entanglement scales as expected: The collapse in Fig. 5c (with $S_{2}/W$ on the $y$ axis) shows that the entropy in the hard (highly entangled) phase is proportional to $W$, and Fig. 5b (with $y$ axis $S_{2}/\log D$) is consistent with convergence to an $S_{2}\propto\log D$ scaling—both are consistent with the maximum possible value $W\log D$ for the entanglement. In the easy (weakly entangled) phase, we observe (not shown) that the data for different $W$ collapses, and decays roughly exponentially with $\lambda$, with a slight $D$-dependence of the decay. Finally, let us point out an interesting transition in the distribution of $S_{2}$ over different random instances for fixed $D$, $W$, and $\lambda$: While in the hard phase, $\exp(S_{2})$ follows a Gaussian distribution (which is natural, given that $S_{2}$ is a logarithmic quantity), in the easy phase $\log(S_{2})$ (!) is Gaussian distributed. The latter in particular implies that the distribution of $S_{2}$ has long tails, and the true average of $S_{2}$ might be somewhat larger; at the same time, it raises the question whether the randomness average is really the most meaningful quantity to study as opposed to, say, the median. Finally, let us point out that the same transition is also observed in other entropies, in particular also those with Rényi index smaller than one, as well as for the von Neumann entropy.
@@ IV.3 Unitary ensemble
This picture offers a few appealing features, and could potentially give rise to alternative sampling-based contractions. It is, however, far from rigorous – neither do we know that the sampling does not have a sign problem, nor is it clear that the fact that most individual realizations of $A$ vs. $S$ tensors have low entanglement is reflected in the entanglement of the overall state. At the same time, it suggests interesting connections to e.g. measurement-based induced phase transitions [22, 23, 24, 25, 26], where probabilistic (or otherwise weak) measurements induce an entanglement transition in the output of a quantum circuit, related to an underlying percolation transition.
@@
Specifically, Fig. 7a shows the entanglement vs. $\lambda$ for an interpolation to a rank-$1$ tensor $S$ with both positive and negative entries. We observe that—unlike in Fig. 5b—(i) there is no sharp transition, but rather a relatively smooth decrease in the entanglement (and, e.g., the tails have much larger entropy), and (ii) the cross-over in the entanglement happens around $\lambda\sim 1$, rather than $\lambda=1/D$ (observe the similarity between the curves for different $D\geq 3$). Indeed, this is consistent with the percolation argument, whose transition depends on the density of $S$, and thus the value of $\lambda$, independent of $D$. Still, it is noteworthy that it is at a value of $\lambda$ on the order of $1$ where the entanglement starts to vanish, which is where we also observed a corresponding behavior for $\Delta f$ (Fig. LABEL:fig:free-energy), even though in the current case, the transition does not seem to depend on positivity, in contrast to the the transition seen in $\Delta f$.
@@ V PEPS expectation values and completely positive tensor networks
(see Fig. 9a), where the double-layer index $x_{\bullet}=(y_{\bullet},z_{\bullet})$, $\bullet=\ell,r,u,d$, has dimension $D^{2}$ for a PEPS with bond dimension $y_{\bullet}=1,\dots,D$. The tensors $A^{(s)}$ of the resulting 2D tensor network thus do not in general have positive entries (unless they are obtained from PEPS with positive entries). They do, however, come with another type of positivity: When interpreted as a map from ket to bra indices, they amount to positive semidefinite operators, i.e., they can be written in the form (13) (i.e., “$C^{\dagger}C$” acting from ket to bra) for some $C^{(s)}$. Equally, the evaluation of physical observables can be decomposed as a sum of two such positive semidefinite objects. This raises the question whether this positivity structure can still ease the computational difficulty of the contraction. In the following, we will address this question through the various tools introduced in the preceding sections.
@@ VI Discussion
Entanglement transitions have been studied in recent years in a variety of contexts. This in particular involves the questions of dynamics and equilibration, such as the MBL transition [34] or entanglement transitions in monitored quantum dynamics, measurement-induced transitions, or other transitions obtained by adding certain non-unitary gates to unitary quantum circuits [22, 23, 24, 25, 26]. Since quantum circuits, including measurements, can be expressed as tensor networks, these systems therefore also provide examples of entanglement transitions in tensor networks. Similarly, entanglement transitions in random tensor networks as a function of the bond dimension $D$ have been investigated previously for random entries with mean 00 [17, 18] (i.e., the subfamily with $\lambda=0$ of the class we investigated), mimicking non-integer $D$ by a suitable random distribution of bond dimensions; there, an entanglement transition between $D=2$ and $D=3$ has been observed, consistent with our findings.
This illustrates that there is a number of different parameters which can drive entanglement transitions in tensor networks. On the one hand, our results add another axis to it, namely, the positivity of the tensor entries. Yet, there is a profound difference in the nature of the sign problem transition which we observe: The root of this transition is in the fundamentally different computational complexity of the underlying computational problem, and the observed entanglement transition is only one possible way in which this complexity transition manifests itself. In fact, in a forthcoming work, we will show that the same transition also appears in an entirely different contraction algorithm which is inspired by Barvinok’s method for approximating the permanent [35, 36], and which makes no reference to entanglement whatsoever [37]. In this context, a key question which remains to be answered is to determine the true point of the complexity transition, that is, to identify a regime for $\lambda$ where contraction is provably hard.
@@
Finally, the connections which we established between the positivity of a tensor network, its entanglement scaling, and the hardness of contracting it might also help to develop improved contraction algorithms. In particular, we observed that PEPS contraction has a favorable entanglement scaling, and we also found that the corresponding two-layer tensor network can be transformed into a positive tensor network by a local TNR-like transformation. In turn, we had observed that positive tensor networks have a favorable entanglement scaling and generally a lower computational complexity. This suggests to use positivity of the entries as one guiding principle for how to perform truncations in TNR or other contraction algorithms. Such a truncation should result in tensor networks which can be much more efficiently contracted in the subsequent steps, and thus might very well lead to overall more accurate contractions when compared to methods which minimize the truncation error made in each step individually.

### v2

@@ III Monte Carlo based contraction
Fig. LABEL:fig:free-energy shows the average $\Delta f$ as a function of $\lambda$ for different bond dimensions $D$, both for Haar-random orthogonal and unitary tensors. [^3] Data has been obtained through transfer operator calculations, applied to a positive vector, on very long cylinders of circumference $W$, measuring the ratio of the change in norm per slice (after discarding the initial head), and taking the average. Computations for different $W$ show a linear scaling in $W$, implying stability of the limit $\lim\tfrac{1}{W}\log\mathcal{C}/\mathcal{C}_{b}$ used in obtaining $\Delta f$. We can identify three different regimes: First, for $\lambda\leq\eta_{0}/D$ (with $\eta_{0}\approx 1$), $\Delta f$ is constant, with a value $\Delta f=\log D+\mathrm{const.}$, as can be seen from the data collapse for all $D$ except $D=2$ in Fig. LABEL:fig:free-energyb (with the constant depending on the choice of the random ensemble). However, once outside this regime, the $\Delta f(\lambda)$ for different $D$ collapse more or less perfectly, with only minor deviations for $D=2$. $\Delta f(\lambda)$ then exhibits two different regimes: For $\eta_{0}/D\leq\lambda\lesssim 1$, we see a behavior $\Delta f\sim-\log\lambda$ (as expected from the collapse in Fig. LABEL:fig:free-energyb and the transition at constant $\lambda D$). For $\lambda\gg 1$, the scaling behavior depends on the chosen distribution—for Haar-orthogonal tensors, $\Delta f$ decays super-exponentially, while for Haar-unitary tensors, $\Delta f\sim 1/4\lambda^{2}$ (Fig. LABEL:fig:free-energyd,e).
@@
From Fig. LABEL:fig:free-energy, we conclude that on the one hand, for very small $\lambda$, the severity of the sign problem depends on the bond dimension of the tensor network, getting worse with larger $D$. Specifically, the $\Delta f=\log D+\mathrm{const.}$ scaling implies that $e^{\Delta f\,N}\sim D^{N}$, which is consistent with the expected worst-case $D^{N}$ scaling (corresponding to exact contraction). Then, for $1/D\lesssim\lambda\lesssim 1$, this transitions into a regime where there is still a sign problem, which now however is independent of the bond dimension. Only for $\lambda\gtrsim 1$, a transition to a rapid decay of $\Delta f$ sets in, meaning that the sign problem rapidly disappears with increasing $\lambda$. This transition is indeed what one would have naively expected, given that around $\lambda\gtrsim 1$, the entries of the tensors start to be predominantly positive, with the probability of negative entries rapidly decreasing as $\lambda$ increases. Concretely, in the orthogonal case, the superexponential decay of $\Delta f(\lambda)$ implies the absence of a sign problem as long as $N$ does not grow superexponentially with $\lambda$, while for the unitary case, the sign problem sets in for $N\sim 4\lambda^{2}$; we discuss the origin of this different behavior in the next paragraph. To give a specific illustration, for the orthogonal case and $\lambda=2$, a $10\times 10$ lattice can be contracted using Monte Carlo sampling without a sign problem, while for $\lambda=3$, one can even go up to a lattice of about $100\times 100$ sites. Note that ultimately, the question whether there is a sign problem depends on the parameter regime we are interested in: If we keep $\lambda$ fixed and increase $N$, any setting will eventually exhibit a sign problem (except for $\lambda\geq D^{2}$ in the orthogonal case where the probability for non-positive entries is strictly zero, see below)—yet, the phenomena of interest are phase transitions in the scaling behavior driven by collective phenomena, rather than transitions which are due to the complete absence of negative entries already locally.
Let us now briefly discuss the origin of the different scaling observed in the $\lambda\gg 1$ regime for Haar-orthogonal vs. Haar-unitary tensors, and its implication on the disappearance of the Monte Carlo sign problem for $\lambda\gtrsim 1$. To this end, we use that the entries of a Haar-random matrix follow a Gaussian distribution to high accuracy. For Haar-orthogonal tensors, the tensors $A_{\bm{x}}^{(s)}$, Eq. (3), thus follow a Gaussian distribution with standard deviation $1$ and mean $\lambda$. The probability for a negative value of $T(\bm{x})$ thus originates from the tail of the Gaussians, which explains their super-exponential decay. On the other hand, for a Haar-unitary distribution, the entries $A_{\bm{x}}^{(s)}=re^{i\phi}$ follow a complex Gaussian distribution with standard deviation $1$ and mean $\lambda$. The fluctation of the complex phase $\phi$ is thus $\sigma(\phi)\sim 1/\lambda$ for $\lambda\gg 1$: This implies that when multiplying $N$ of those numbers, as is the case for the values $T(\bm{x})$ of the tensor network for a fixed setting of the indices, the overall phase gets randomized when $\sqrt{N}\sigma(\phi)\sim 1$ and thus, following Eq. (4), $\Delta f\sim 1/N\sim 1/\lambda^{2}$.
For comparison, we have also considered tensor networks with real or complex Gaussian entries (shown in Fig. LABEL:fig:free-energy with solid semi-transparent markers), which show the same behavior to very high accuracy. One difference, however, is the behavior for large $\lambda$ in the real case, as for Haar-orthogonal tensors, $\Delta f\equiv 0$ as soon as $\lambda\geq D^{2}$ (as all entries of a unitary have magnitude $\leq 1$), while for Gaussian tensors, $\Delta f$ will always have some finite non-zero value due to the tail of the Gaussian, and thus, the decay for large $\lambda$ must be more rapid; this can indeed be be seen in Fig. LABEL:fig:free-energyd,e for the $D=2$ data.
@@ IV Entanglement based contraction
Figure 4: Contraction of a 2D tensor network. One proceeds by blocking columns (blue box in panel (a)). The resulting block $\psi$ is then approximated by a 1D tensor network, i.e., an MPS, with bond dimension $\chi$ (panel (b)). In order to keep the truncation error in all iterations small, it is necessary that the correlations between the upper and lower part of $\psi$ are bounded, i.e., the singular value decomposition in the $A$ vs. $B$ partition has a small tail (shown in panel (c) across the central cut, where we expect most correlations.) This weight can be assessed by whether the entropy across the cut scales with the width $W$. By choosing $H/2$ sufficiently larger than $W$, we ensure that solely $W$ bounds the possible correlations. Concretely, we chose $H/2=2W$: This value is sufficiently larger than the value $H/2\approx W$ where the entanglement starts to saturate, and we have checked numerically that further increasing $H$ does not lead to any change within error bars.
@@ IV.1 Numerical results
Leaving the surprising transition point aside, the entanglement scales as expected: The collapse in Fig. 5c (with $S_{2}/W$ on the $y$ axis) shows that the entropy in the hard (highly entangled) phase is proportional to $W$, and Fig. 5b (with $y$ axis $S_{2}/\log D$) is consistent with convergence to an $S_{2}\propto\log D$ scaling—both are consistent with the maximum possible value $W\log D$ for the entanglement. In the easy (weakly entangled) phase, we observe (not shown) that the data for different $W$ collapses, and decays roughly exponentially with $\lambda$, with a slight $D$-dependence of the decay. Finally, let us point out that the same transition is also observed in other entropies, in particular also those with Rényi index smaller than one, as well as for the von Neumann entropy.
@@ IV.3 Unitary ensemble
This picture offers a few appealing features, and could potentially give rise to alternative sampling-based contractions. It is, however, far from rigorous – neither do we know that the sampling does not have a sign problem, nor is it clear that the fact that most individual realizations of $A$ vs. $S$ tensors have low entanglement is reflected in the entanglement of the overall state. At the same time, it suggests interesting connections to e.g. measurement-based induced phase transitions [22, 23, 24, 25, 26, 27], where probabilistic (or otherwise weak) measurements induce an entanglement transition in the output of a quantum circuit, related to an underlying percolation transition.
@@
Specifically, Fig. 7a shows the entanglement vs. $\lambda$ for an interpolation to a rank-$1$ tensor $S$ with both positive and negative entries. We observe that—unlike in Fig. 5b—(i) there is no sharp transition, but rather a relatively smooth decrease in the entanglement (and, e.g., the tails have much larger entropy), and (ii) the cross-over in the entanglement happens around $\lambda\sim 1$, rather than $\lambda=1/D$ (observe the similarity between the curves for different $D\geq 3$). Indeed, this is consistent with the percolation argument, whose transition depends on the density of $S$, and thus the value of $\lambda$, independent of $D$ (and indeed, such a percolation transition at fixed finite probability $p_{c}$—the percolation threshold—is what is observed also in measurement-induced phase transitions [22, 23, 24, 25, 26, 27]). Still, it is noteworthy that it is at a value of $\lambda$ on the order of $1$ where the entanglement starts to vanish, which is where we also observed a corresponding behavior for $\Delta f$ (Fig. LABEL:fig:free-energy), even though in the current case, the transition does not seem to depend on positivity, in contrast to the the transition seen in $\Delta f$.
@@ V PEPS expectation values and completely positive tensor networks
(see Fig. 9a), where the double-layer index $x_{\bullet}=(y_{\bullet},z_{\bullet})$, $\bullet=\ell,r,u,d$, has dimension $D^{2}$ for a PEPS with bond dimension $y_{\bullet}=1,\dots,D$. The tensors $A^{(s)}$ of the resulting 2D tensor network thus do not in general have positive entries (unless they are obtained from PEPS with positive entries). They do, however, come with another type of positivity: When interpreted as a map from ket to bra indices, they amount to positive semidefinite operators, i.e., they can be written in the form (13) (i.e., “$C^{\dagger}C$” acting from ket to bra) for some $C^{(s)}$. This is also true for computing expectation values of of physical observables, since any observable can be expressed as the difference of two positive semidefinite operators (and even for string operators, those positive semidefinite operators can be decomposed in terms of a double-layer tensor network). This raises the question whether this positivity structure can still ease the computational difficulty of the contraction. In the following, we will address this question through the various tools introduced in the preceding sections.
@@ VI.1 Summary
VI.2 Context
Entanglement transitions have been studied in recent years in a variety of contexts. This in particular involves the questions of dynamics and equilibration, such as the MBL transition [35] or entanglement transitions in monitored quantum dynamics, measurement-induced transitions, or other transitions obtained by adding certain non-unitary gates to unitary quantum circuits [22, 23, 24, 25, 26, 27]. Since quantum circuits, including measurements, can be expressed as tensor networks, these systems therefore also provide examples of entanglement transitions in tensor networks. Similarly, entanglement transitions in random tensor networks as a function of the bond dimension $D$ have been investigated previously for random entries with mean 00 [17, 18] (i.e., the subfamily with $\lambda=0$ of the class we investigated), mimicking non-integer $D$ by a suitable random distribution of bond dimensions; there, an entanglement transition between $D=2$ and $D=3$ has been observed, consistent with our findings.
This illustrates that there is a number of different parameters which can drive entanglement transitions in tensor networks. On the one hand, our results add another axis to it, namely, the positivity of the tensor entries. Yet, there is a profound difference in the nature of the sign problem transition which we observe: The root of this transition is in the fundamentally different computational complexity of the underlying computational problem, and the observed entanglement transition is only one possible way in which this complexity transition manifests itself. In fact, in a forthcoming work [38], we will show that the same transition also appears in an entirely different contraction algorithm which is inspired by Barvinok’s method for approximating the permanent [36, 37], which makes no reference to entanglement whatsoever, and which can be rigorously proven to work. In this context, a key question which remains to be answered is to determine the true point of the complexity transition, that is, to identify a regime for $\lambda$ where contraction is provably hard.
@@ VI.2 Context
VI.3 Practical implications
Finally, let us discuss how the results of this work, which focused on the study of random tensor networks, can be relevant for practical settings, in particular for the simulation of physical many-body problems based on tensor network methods.
A first point to note is that random tensor networks with positive entries, as considered in this work, naturally appear as partition functions of random classical statistical mechanics models, which are widely studied in the context of disordered many-body systems. Thus, the results of the analysis of random tensor networks which we obtained in this work directly carry over to random classical statistical mechanics models.
Next, let us turn towards the possible implications of our results for numerical simulation algorithms for quantum many-body systems based on tensor networks, in particular variational PEPS simulations. A key insight of our work is that positive tensor networks—and, in fact, already tensor networks which are biased towards positive entries—possess a much lower amount of entanglement and can therefore be more efficiently contracted using commonly used contraction algorithms. This suggests that positivity of tensor entries can serve as a guide in choosing the most suitable gauge when contracting tensor networks in PEPS simulations: Gauge choices can have a significant effect on the efficiency of contracting a tensor network, and at the same time, choosing the right gauge is a challenging problem in many cases.
Figure 10: Contraction of a translational invariant infinite PEPS obtained by optimizing a highly frustrated topological spin liquid (square lattice $J_{1}$-$J_{2}$ model at $J_{2}/J_{1}=0.5$) which exhibits a strong sign problem. The plot on the left shows the truncation error $\delta$ (sum of discarded Schmidt coefficients) for the left and right transfer matrix fixed point as a function of the truncation dimension $\chi$ (cf. Fig. 4b), while the table on the top right shows the $2$-Renyi entanglement entropies $S_{2}$ for a bipartition of those fixed points. We compare data obtained for the original iPEPS tensor returned by a variational minimization (blue) with data obtained for tensors where we have optimized over gauge choices $X$ and $Y$ as shown on the bottom right. On the one hand, we have optimized $X$ and $Y$ such as to maximize the hermiticity of the transfer matrix (green), and on the other hand, we have additionally optimized $X$ and $Y$ for positivity of the tensor entries (red). We see that both optimizations lead to a situation where the truncation error and entanglement entropy for the two fixed points become symmetric. However, while optimizing for hermiticity leads to an increased truncation error and entanglement entropy for one of the fixed points (such that the optimized, symmetric values are in between the original values), optimizing for positivity allows to decrease the truncation error for both fixed points simultaneously and thus gives rise to a more efficient contraction of the tensor network.
One possibility how positivity of tensor entries can guide such a gauge choice is as follows (for clarity, we focus on the practically most relevant case of a translational invariant infinite tensor network). We start from the two-layer tensor network obtained when computing expectation values, which is made up from individual two-layer ket-bra tensors $A$ just as the one shown in Fig. 9a. The goal is to contract this tensor network on an infinite system using infinite boundary MPS (iMPS) (i.e., the translational invariant version of Fig. 4b) from both left and right; from the iMPS fixed points, we can then compute physical quantities in the given state. The value of this tensor network is invariant under gauge transformations $X$, $X^{-1}$ and $Y$, $Y^{-1}$ applied to the horizontal and vertical indices, respectively, as shown in Fig. 10. We can now choose the gauges $X$ and $Y$ such as to minimize the amount of non-positive tensor entries, e.g. by maximizing $\big{|}\sum_{\bm{x}}A_{\bm{x}}\big{|}/\sum_{\bm{x}}\big{|}A_{\bm{x}}\big{|}$, where the sum runs over all entries of the double-layer tensor $A_{\bm{x}}$.
To test the effect of such a gauge fixing in a practically relevant scenario, we have used a tensor network state (infinite PEPS) which has been variationally optimized for the ground state of the spin-$\tfrac{1}{2}$ $J_{1}$-$J_{2}$-model on the square lattice for $J_{2}/J_{1}=0.5$ (i.e., a Heisenberg antiferromagnet with coupling $J_{1}$ along the edges and coupling $J_{2}$ along the diagonal). At this value of $J_{2}/J_{1}$, the model is a highly frustrated topological spin liquid, and exhibits a strong sign problem in quantum Monte Carlo. For this tensor, we have then studied how a gauge fixing procedure guided by positivity of the tensor entries affects the amount of entanglement in the boundary iMPS and thus the hardness of contraction. The results are reported in Figure 10 and the figure caption, where we compare the contraction of the network constructed using the original double-layer tensor, the tensor where the gauge has been chosen such as to maximize the hermiticity of the the transfer operator, and the tensor where the gauge has been chosen such as to maximize the positivity of the entries while at the same time favoring hermiticity. We find that choosing a gauge which maximizes positivity leads to a significant decrease both of the entanglement entropy and the truncation error of the boundary iMPS in the contraction. This demonstrates that the observations made in our work about the relation of entanglement and positivity also apply in realistic, translational invariant settings involving systems with a severe sign problem, and that these insights can thus be used to improve the accuracy of practical contraction algorithms.
Another way in which the findings reported in this work could be used to improve contraction algorithms in practical settings builds on the observation made in Sec. V.3, where we showed that in a certain regime, the two-layer tensor network which appears in PEPS contractions can be transformed into a positive tensor network by a local TNR-like transformation. In a realistic regime, we wouldn’t expect such a strictly positive TNR decomposition to typically exist. However, as we observed, a relatvely small positive bias can already be sufficient to significantly reduce the complexity in tensor network contraction. This suggests to construct a TNR-type transformation whose aim is to give rise to a positively biased tensor network after one TNR step, Fig. 9c,d—this could be achieved, e.g., by replacing the separable decomposition in Eq. (16) by an approximately separable decomposition, rather than the singular value decomposition typically used in TNR. Such a positively biased TNR procedure could then, e.g., be used as an initial pre-processing step in tensor network contraction, thereby lowering the amount of entanglement in the tensor network, such that the resulting tensor network could subsequently be contracted much more efficiently.
Lastly, let us give one more argument which illustrates why one should expect findings for random to be relevant also for tensors which appear in practical simulations. It is based on the strong concentration effects observed around the average, an effect which gets very rapidly stronger with bond dimension [28]. Thus, a random instance will behave exactly like the average, with a probability extremely close to one. If we now start from a random PEPS and optimize it to be the ground state of a given Hamiltonian, we will end up with a PEPS in close vicinity of the true ground state, but which we otherwise (i.e., within this vicinity) expect to be completely random. The strong concentration effects thus imply that with probability rapidly approaching one (for increasing bond dimension, which is the regime we aim for in variational algorithms), such a variational PEPS optimization will end up in a PEPS wavefunction which shares the same properties (in particular regarding the efficient computability of expectation values) as the random instances studied in our work.
Finally, the results of our work strengthen the case of using entanglement-based algorithms for tensor network contraction, such as the commonly used methods based on boundary MPS or the corner transfer matrix, over Monte Carlo based approaches. Firstly, we have observed that in order to reduce the difficulty of entanglement-based approaches, a significantly lower bias towards positive entries is required as compared to Monte Carlo sampling; in fact, the reduction of difficulty happens earlier for instances with higher bond dimension (which one would expect to be more difficult, and which have a more severe sign problem in Monte Carlo). In addition, for Monte Carlo based approaches at any fixed positive bias $\lambda$, the sign problem will eventually kick in as the system size $N$ is increased, even though this might happen rather late for larger $\lambda$. In contrast to this, for entanglement based approaches, the amount of entanglement no longer scales with $N$ once the positive bias exceeds $\lambda\approx 1/D$, and thus, arbitrarily large systems can be simulated at a cost which only grows linearly with $N$ or, in the case of infinite systems, becomes system size independent. Therefore, our findings further strengthen the case to favor entanglement based methods over Monte Carlo sampling in tensor network contraction.